OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = x2 - 2x + 1, what is f(g(x)) ? 4x2 - 8x + 5 4x2 + 2x + 2 x2 + 2x + 2 x2 - 2x + 2 which one?

OpenStudy (anonymous):

f(g(x))= 4(g(x))+1= 4(x2-2x+1)+1 = 4x2 - 8x +4 +1 = 4x2 -8x +5